### v1

@@ 2.5 The Communication Process: Downlink Data Transmission
We assume that each user may be served by a certain subset of APs and BSs, and denote by $a_{k,m}$ and $b_{k,\ell}$ the binary variates encoding this information. In particular, $a_{k,m}=1$ ($b_{k,\ell}=1$) of the $k$-th user is served by the $m$-th AP ($\ell$-th BS), otherwise it is zero. We also assume that the coefficients $a_{k,m}$ ($b_{k,\ell}$) are arranged into the matrix $\mathbf{A}$ ($\mathbf{B}$) of size $K\times M$ ($K\times L$). The APs and the BSs treat the channel estimates as the true channels and perform beamforming on the downlink. We denote by $\mathbf{w}_{k,m}^{(a)}$ ($\mathbf{w}_{k,\ell}^{(b)}$) the unit-norm beamforming vector of dimension $N_{\rm AP}$ ($N_{\rm BS}$) used at the $m$-th AP ($\ell$-th BS) for the downlink transmission to the $k$-th user, and by $\eta_{k,m}^{(a)}$ ($\eta_{k,\ell}^{(b)}$) the positive scalar coefficient controlling the power transmitted by the $m$-th AP ($\ell$-th BS) to the $k$-th user. Moreover, Letting $P^{(a),{\rm DL}}_{m}$ $\left(P^{(b),{\rm DL}}_{\ell}\right)$ denote the maximum transmitted power by the $m$-th AP ($\ell$-th BS), the normalized transmit powers must satisfy the constraints
$\sum_{k=1}^{K}{\eta_{k,m}^{(a)}}\leq P^{(a),{\rm DL}}_{m}\,,\quad\mbox{and}\quad\displaystyle\sum_{k=1}^{K}{\eta_{k,\ell}^{(b)}}\leq P^{(b),{\rm DL}}_{\ell}\;.$ (7)
Based on the above assumptions, following the usual transceiver signal processing steps in downlink cell-free user-centric systems, it is straightforward to show that the $k$-th user downlink SINR, $\gamma_{k}^{\rm DL}$ say, conditioned on the fast fading channel realizations and on the channel vector estimates is written as
$\!\!\gamma_{k}^{\rm DL}\!=\!\frac{\displaystyle\left|\displaystyle\sum_{m=1}^{M}\!\!a_{k,m}\displaystyle\sqrt{\eta_{k,m}^{(a)}}\mathbf{g}_{k,m}^{H}\mathbf{w}_{k,m}^{(a)}\!+\!\displaystyle\sum_{\ell=1}^{L}b_{k,\ell}\displaystyle\sqrt{\eta_{k,\ell}^{(b)}}\mathbf{h}_{k,\ell}^{H}\mathbf{w}_{k,\ell}^{(b)}\right|^{2}}{\displaystyle\sum_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{K}\!\left|\displaystyle\!\sum_{m=1}^{M}\!\!a_{j,m}\displaystyle\!\sqrt{\!\eta_{j,m}^{(a)}}\mathbf{g}_{k,m}^{H}\!\mathbf{w}_{j,m}^{(a)}\!\!+\!\!\displaystyle\sum_{\ell=1}^{L}\!b_{j,\ell}\!\displaystyle\sqrt{\!\eta_{j,\ell}^{(b)}}\mathbf{h}_{j,\ell}^{H}\!\mathbf{w}_{j,\ell}^{(b)}\right|^{2}\!\!+\!\sigma^{2}_{z}}\,,$ (8)
with $\sigma^{2}_{z}$ the AWGN noise at the $k$-th MS.
@@ 3 Downlink spectral efficiency bounds
The performance measure that is considered in this work is the per-user downlink rate, that can be obtained by multiplying the downlink per-user spectral efficiency (SE) by the signal bandwidth. Unfortunately, the exact closed-form expression of the downlink SE is not available, and suitable bounds are to be adopted. Assuming a Gaussian approximation for the interference, Gaussian distributed very long codewords, and perfect channel state information (PCSI), the ergodic SE for the $k$-th user is written as
@@
Under the practical case of imperfect channel state information (ICSI), the channel coefficients are estimated in the uplink, and the estimated values are used to compute the beamformers. Let us thus denote by $\widehat{\gamma}_{k}^{\rm DL}$ the $k$-th user SINR expression in (8), where the beamformers $\mathbf{w}_{k,m}^{(a)}$ and $\mathbf{w}_{k,\ell}^{(b)}$, for all values of $k,m$ and $\ell$, are computed based on the available channel estimates. It can be then shown that, for the ICSI case, an upper bound to the achievable SE for the user $k$ is written as
@@
Besides the upper-bound (10), some lower-bound expressions for the per-user downlink SE. These include the Use-and-then-Forget (UatF) bound [28], and the Caire bound in [29]. The former bound had been originally conceived for massive MIMO systems, and was extended to cell-free massive MIMO systems in [30]. This bound, however, has been shown to be somehow loose for cell-free massive MIMO systems, since it works well in the presence of channel hardening, a phenomenon that takes place in the massive MIMO scenario where BSs are equipped with a large number of co-located antennas. Moreover, a closed-for expression of the beamformer is available only for a limited number of beamformers type. To circumvent these drawbacks, in [29] two alternative lower bounds were proposed; these ones, however, are tight when the channel coherence block length is large with respect to the number of users, and in the case where interference is nearly entirely eliminated by zero-forcing beamforming. Clearly, these assumption are hardly verified in the practically relevant case of an overloaded network with a large number of active users and a non-negligible level of interference. Accordingly, in this paper, we will use the upper boud (10) as our key performance indicator. Indeed, the bound tends to become tight when the channel estimation error decreases; moreover, although showing SE values not practically achievable, it is still useful when comparing different network deployments, network cooperation levels, and different implementations of the beamformers and of the power control rules.
@@ 4 System design
In this section, we discuss the considered UE-AP/BS association rules, the adopted beamformers and power control strategies.
@@ 4.1 Basic UE-AP/BS association rules
The optimal association between each UE and the serving APs and BSs descends from the adopted optimization criterion, and is generally an NP-hard combinatorial problem. Accordingly, heuristic association rules are considered in practice based on the large scale fading coefficients.
For the MC-mMIMO scenario (i.e., multicell massive MIMO with no BS cooperation), each UE is simply associated to the BS with the largest value of large scale fading coefficient. Otherwise stated, letting $\ell^{*}_{k}=\arg\max_{\ell}\rho_{k,\ell}$, we have that $b_{k,\ell^{*}_{k}}=1$ and $b_{k,\ell}=0$, for all $\ell\neq\ell^{*}_{k}$, and the matrix $\mathbf{A}$ consists of all zeros.
For the HET-noCoop scenario, each UE can connect either to a certain number, say $\widetilde{N}_{\rm BS}$, of BSs or to a certain number, say $\widetilde{N}_{\rm AP}$, of APs. The association rule has to take into account the strength of the large scale fading coefficients and the number of antennas at the APs and BSs. A suitable association rule can be obtained as follows:
@@
    then the $k$-th UE is to be associated to the APs with index $O^{\beta}_{k}(1),O^{\beta}_{k}(2),\ldots,O^{\beta}_{k}(\widetilde{N}_{\rm AP})$. Accordingly, coefficients $a_{k,O^{\beta}_{k}(1)},\ldots,a_{k,O^{\beta}_{k}(\widetilde{N}_{\rm AP})}$ should be set equal to 1, and all the other entries of the matrix $\mathbf{A}$ and the entire matrix $\mathbf{B}$ should be set to zero.
@@
    If, conversely, (11) does not hold, then the $k$-th UE is to be associated to the BSs with index $O^{\rho}_{k}(1),O^{\rho}_{k}(2),\ldots,O^{\rho}_{k}(\widetilde{N}_{\rm BS})$. Accordingly, coefficients $b_{k,O^{\rho}_{k}(1)},\ldots,b_{k,O^{\rho}_{k}(\widetilde{N}_{\rm BS})}$ should be set equal to 1, and all the other entries of the matrix $\mathbf{B}$ and the entire matrix $\mathbf{A}$ should be set to zero.
For the HET-noCoop scenario, where each UE can connect to either one BS or to one AP, the UE-AP/BS association rule is a special case of the one outlined for the previous HORIZONTAL-Coop scenario, obtained by setting $\widetilde{N}_{\rm BS}=\widetilde{N}_{\rm AP}=1$.
Finally, in the FULL-Coop scenario, each UE is cooperatively served by $\widetilde{N}_{\rm AP}$ APs and by $\widetilde{N}_{\rm BS}$ BSs. Also for this scenario the large scale fading coefficients are considered for selecting the serving APs and BSs. In particular, using again the sorting operators $O_{k}^{\beta}$ and $O_{k}^{\rho}$, we have that the association variables $a_{k,O^{\beta}_{k}(1)},\ldots,a_{k,O^{\beta}_{k}(\widetilde{N}_{\rm AP})}$ and $b_{k,O^{\rho}_{k}(1)},\ldots,b_{k,O^{\rho}_{k}(\widetilde{N}_{\rm BS})}$ are set equal to one, while the remaining entries of the matrices $\mathbf{A}$ and $\mathbf{B}$ are set to zero.
The above described association rule are, as already highlighted, simple and heuristic. Alternative and more sophisticated rules can be thus conceived. In particular, the association could take into account not only the channel strenght, as the described policies do, but also the interference level. Moreover, load balancing constraint could be also introduced, so as to prevent the case in which a single AP or BS is connected to a too large number of devices. The final part of this paper will provide results along these lines.
@@ 4.2 Downlink beamforming
Once the UE-AP/BS association rule has been implemented, the set $\mathcal{K}_{m}$ ($\mathcal{J}_{\ell}$) can be defined for the generic $m$-th AP ($\ell$-th BS). In particular, $\mathcal{K}_{m}$ is the set of the users served by the $m$-th AP, i.e., $\mathcal{K}_{m}=\{k\,:\,a_{k,m}=1\}$. Similarly, $\mathcal{J}_{\ell}$ is the the set of the users served by the $\ell$-th BS, i.e., $\mathcal{J}_{\ell}=\{k\,:\,b_{k,\ell}=1\}$.
The following three downlink beamformers will be considered in this paper.
@@
Partial zero-forcing (PZF): In this case, we assume that each AP performs a local PZF downlink precoding, i.e. it nulls its interference contribution towards some selected users. Consider the $m$-th AP and assume that $k\in\mathcal{K}_{m}$. Denote by $N_{\rm PZF}^{\rm AP}$ the number of users to be protected by the interference generated by each AP [^1] Recall that $N_{\rm PZF}^{\rm AP}$ must not exceed the number of antennas available at the AP. when transmitting to UE $k$. Denote by $j_{k,m}(1),\ldots,j_{k,m}(N_{\rm PZF}^{\rm AP})$ the index of the users to be protected [^2] A customary choice may be to select the users, other than the $k$-th, with the strongest large scale fading coefficient towards the $m$-th AP.; the PZF precoder to be used at $m$-th AP to transmit to the $k$-th user is computed as follows:
@@
    Let $\widetilde{\mathbf{I}}_{k,m}=\mbox{orth}\left[\mathbf{I}_{k,m}\right]$ be a matrix containing on its columns an orthonormal basis for the column span of $\mathbf{I}_{k,m}$.
@@
A similar procedure is used in order to determine the beamformer at the $\ell$-th BS for transmitting at UE $k$. Denote by $N_{\rm PZF}^{\rm BS}<N_{\rm BS}$ the number of users to be protected by the interference generated by the BS $\ell$ when transmitting to UE $k$. Denote by $j_{k,\ell}(1),\ldots,j_{k,\ell}(N_{\rm PZF}^{\rm BS})$ the index of the users to be protected [^3] Again, these can be the users, other than the $k$-th, with the strongest large scale fading coefficient towards BS $\ell$.; the PZF precoder to be used at the $\ell$-th BS to transmit to $k$-th user thus requires the formation of the interference matrix $\mathbf{I}_{k,\ell}=[\widehat{\mathbf{h}}_{j_{k,\ell}(1),\ell},\ldots,\widehat{\mathbf{h}}_{j_{k,\ell}(N_{\rm PZF}^{\rm BS}),\ell}]$, containing the estimated channel of the users to be protected, the computation of an orthonormal basis $\widetilde{\mathbf{I}}_{k,\ell}=\mbox{orth}\left[\mathbf{I}_{k,\ell}\right]$, and of the beamformer: Define the PZF
@@
Linear minimum mean squares error (LMMSE): The LMMSE beamforming vector is obtained exploiting the duality principle in massive MIMO systems and minimizing the conditional mean-squares error (MSE) in the uplink decoding[31]. Assuming the LMMSE channel estimation discussed in Section 2.3, the LMMSE beamforming vectors computed at the $m$-th AP and at the $\ell$-th BS for the $\mathcal{K}_{m}(i)$-th user and the $\mathcal{J}_{\ell}(i)$-th user are reported at the top of next page, respectively.
$\mathbf{w}_{\mathcal{K}_{m}(i),m}^{(a),{\rm LMMSE}}=\frac{\left\{\displaystyle\sum_{j\in\mathcal{K}_{m}}\eta_{j}\left[\widehat{\mathbf{g}}_{j,m}\widehat{\mathbf{g}}_{j,m}^{H}+(\beta_{j,m}-\gamma_{j,m})\mathbf{I}_{N_{\rm AP}}\right]+\sigma^{2}_{z}\mathbf{I}_{N_{\rm AP}}\right\}^{-1}\widehat{\mathbf{g}}_{\mathcal{K}_{m}(i),m}}{\left\lVert\left\{\displaystyle\sum_{j\in\mathcal{K}_{m}}\eta_{j}\left[\widehat{\mathbf{g}}_{j,m}\widehat{\mathbf{g}}_{j,m}^{H}+(\beta_{j,m}-\gamma_{j,m})\mathbf{I}_{N_{\rm AP}}\right]+\sigma^{2}_{z}\mathbf{I}_{N_{\rm AP}}\right\}^{-1}\widehat{\mathbf{g}}_{\mathcal{K}_{m}(i),m}\right\rVert}\,,$ (13)
$\mathbf{w}_{\mathcal{J}_{\ell}(i),\ell}^{(b),{\rm LMMSE}}=\frac{\left\{\displaystyle\sum_{j\in\mathcal{J}_{\ell}}\eta_{j}\left[\widehat{\mathbf{h}}_{j,\ell}\widehat{\mathbf{h}}_{j,\ell}^{H}+(\rho_{j,\ell}-\delta_{j,\ell})\mathbf{I}_{N_{\rm BS}}\right]+\sigma^{2}_{z}\mathbf{I}_{N_{\rm BS}}\right\}^{-1}\widehat{\mathbf{h}}_{\mathcal{J}_{\ell}(i),\ell}}{\left\lVert\left\{\displaystyle\sum_{j\in\mathcal{J}_{\ell}}\eta_{j}\left[\widehat{\mathbf{h}}_{j,\ell}\widehat{\mathbf{h}}_{j,\ell}^{H}+(\rho_{j,\ell}-\delta_{j,\ell})\mathbf{I}_{N_{\rm BS}}\right]+\sigma^{2}_{z}\mathbf{I}_{N_{\rm BS}}\right\}^{-1}\widehat{\mathbf{h}}_{\mathcal{J}_{\ell}(i),\ell}\right\rVert}\,,$ (14)

### v2

@@ 2.5 The Communication Process: Downlink Data Transmission
We assume that each user may be served by a certain subset of APs and BSs, and denote by $a_{k,m}$ and $b_{k,\ell}$ the binary variates encoding this information. In particular, $a_{k,m}=1$ ($b_{k,\ell}=1$) if the $k$-th user is served by the $m$-th AP ($\ell$-th BS), otherwise it is zero. We also assume that the coefficients $a_{k,m}$ ($b_{k,\ell}$) are arranged in the matrix $\mathbf{A}$ ($\mathbf{B}$) of size $K\times M$ ($K\times L$). The APs and the BSs treat the channel estimates as true channels and perform beamforming on the downlink. We denote by $\mathbf{w}_{k,m}^{(a)}$ ($\mathbf{w}_{k,\ell}^{(b)}$) the unit norm beamforming vector of dimension $N_{\rm AP}$ ($N_{\rm BS}$) used at the $m$-th AP ($\ell$-th BS) for the downlink transmission to the $k$-th user, and by $\eta_{k,m}^{(a)}$ ($\eta_{k,\ell}^{(b)}$) the positive scalar coefficient controlling the power transmitted by the $m$-th AP ($\ell$-th BS) to the $k$-th user. Moreover, letting $P_{m,{\rm max}}^{(a)}$ $\left(P_{\ell,{\rm max}}^{(b)}\right)$ denote the maximum transmitted power by the $m$-th AP ($\ell$-th BS), the normalized transmit powers must satisfy the constraints
$\sum_{k=1}^{K}{\eta_{k,m}^{(a)}}\leq P_{m,{\rm max}}^{(a)}\,,\quad\mbox{and}\quad\displaystyle\sum_{k=1}^{K}{\eta_{k,\ell}^{(b)}}\leq P_{\ell,{\rm max}}^{(b)}\;.$ (9)
Based on the above assumptions, in downlink cell-free user-centric systems, each user receives phase-aligned contributions from the serving APs and BSs. In particular, in a generic symbol interval, the $k$-th user receives the signal expressed in Eq. (10) at the top of the next page, with $x_{k}^{\rm DL}$ the unit power data symbols intended for the $k$-th user, and $z_{k}$ the additive white Gaussian noise (AWGN) with entries ${\cal CN}(0,\sigma^{2}_{z})$.
$\begin{array}[]{llll}r_{k}^{\rm DL}&=\left(\displaystyle\sum_{m=1}^{M}a_{k,m}\displaystyle\sqrt{\eta_{k,m}^{(a)}}\mathbf{g}_{k,m}^{H}\mathbf{w}_{k,m}^{(a)}+\displaystyle\sum_{\ell=1}^{L}b_{k,\ell}\displaystyle\sqrt{\eta_{k,\ell}^{(b)}}\mathbf{h}_{k,\ell}^{H}\mathbf{w}_{k,\ell}^{(b)}\right)x_{k}^{\rm DL}\\ &\quad+\displaystyle\sum_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{K}\left(\displaystyle\sum_{m=1}^{M}a_{j,m}\displaystyle\sqrt{\eta_{j,m}^{(a)}}\mathbf{g}_{k,m}^{H}\mathbf{w}_{j,m}^{(a)}+\displaystyle\sum_{\ell=1}^{L}b_{j,\ell}\displaystyle\sqrt{\eta_{j,\ell}^{(b)}}\mathbf{h}_{k,\ell}^{H}\mathbf{w}_{j,\ell}^{(b)}\right)x_{j}^{\rm DL}+z_{k}\;,\end{array}$ (10)
Given (10), the $k$-th user downlink SINR, $\gamma_{k}^{\rm DL}$ say, conditioned on the fast fading channel realizations and on the channel vector estimates can be easily shown to be written as
$\!\!\gamma_{k}^{\rm DL}\!=\!\frac{\displaystyle\left|\displaystyle\sum_{m=1}^{M}\!\!a_{k,m}\displaystyle\sqrt{\eta_{k,m}^{(a)}}\mathbf{g}_{k,m}^{H}\mathbf{w}_{k,m}^{(a)}\!+\!\displaystyle\sum_{\ell=1}^{L}b_{k,\ell}\displaystyle\sqrt{\eta_{k,\ell}^{(b)}}\mathbf{h}_{k,\ell}^{H}\mathbf{w}_{k,\ell}^{(b)}\right|^{2}}{\displaystyle\sum_{\begin{subarray}{c}j=1\\ j\neq k\end{subarray}}^{K}\!\left|\displaystyle\!\sum_{m=1}^{M}\!\!a_{j,m}\displaystyle\!\sqrt{\!\eta_{j,m}^{(a)}}\mathbf{g}_{k,m}^{H}\!\mathbf{w}_{j,m}^{(a)}\!\!+\!\!\displaystyle\sum_{\ell=1}^{L}\!b_{j,\ell}\!\displaystyle\sqrt{\!\eta_{j,\ell}^{(b)}}\mathbf{h}_{j,\ell}^{H}\!\mathbf{w}_{j,\ell}^{(b)}\right|^{2}\!\!+\!\sigma^{2}_{z}}\,.$ (11)
@@ 3 Downlink spectral efficiency bounds
The performance measure that is considered in this work is the per-user downlink rate, which can be obtained by multiplying the downlink per-user spectral efficiency (SE) by the signal bandwidth. Unfortunately, the exact closed-form expression of the downlink SE is not available, and suitable bounds are to be adopted. Assuming a Gaussian approximation for the interference, Gaussian distributed very long codewords, and perfect channel state information (PCSI), the ergodic SE for the $k$-th user is written as
@@
In the practical case of imperfect channel state information (ICSI), the channel coefficients are estimated in the uplink, and the estimated values are used to compute the beamformers. Let us thus denote by $\widehat{\gamma}_{k}^{\rm DL}$ the $k$-th user SINR expression in (11), where the beamformers $\mathbf{w}_{k,m}^{(a)}$ and $\mathbf{w}_{k,\ell}^{(b)}$, for all values of $k,m$ and $\ell$, are computed based on the available channel estimates. It can be then shown that, for the ICSI case, an upper bound to the achievable SE for the user $k$ is written as
@@
In addition to the upper bound (13), there are some lower bound expressions for the per user downlink SE. These include the Use-and-then-Forget (UatF) bound [32], and the Caire bound in [33]. The former bound had been originally conceived for massive MIMO systems, and was extended to cell-free massive MIMO systems in [7]. However, this bound has been shown to be somehow loose for cell-free massive MIMO systems, since it works well in the presence of channel hardening, a phenomenon that takes place in the massive MIMO scenario where BSs are equipped with a large number of co-located antennas. Moreover, a closed-form expression of the beamformer is available only for a limited number of beamformer types. To avoid these drawbacks, in [33] two alternative lower bounds were proposed; these, however, are tight when the channel coherence block length is large with respect to the number of users, and in the case where interference is nearly entirely eliminated by zero-forcing beamforming. Clearly, these assumptions are hardly verified in the practically relevant case of an overloaded network with a large number of active users and a non-negligible level of interference. Consequently, in this paper, we will use the upper boud (13) as our key performance indicator. In fact, the bound tends to become tight when the channel estimation error decreases; moreover, although showing SE values not practically achievable, it is still useful when comparing different network deployments, network cooperation levels, and different implementations of beamformers and power control rules.
@@ 4 System design
In this section, we discuss the UE-AP/BS association rules considered, the adopted beamformers, and power control strategies.
@@ 4.1 Basic UE-AP/BS association rules
The optimal association between each UE and the serving APs and BSs descends from the adopted optimization criterion and is generally a NP-hard combinatorial problem. Accordingly, heuristic association rules are considered in practice based on the large-scale fading coefficients.
For the MC-mMIMO scenario (i.e., multicell massive MIMO without BS cooperation), each UE is simply associated to the BS with the highest value of the large-scale fading coefficient. Otherwise stated, letting $\ell^{*}_{k}=\arg\max_{\ell}\rho_{k,\ell}$, we have $b_{k,\ell^{*}_{k}}=1$ and $b_{k,\ell}=0$, for all $\ell\neq\ell^{*}_{k}$, and the matrix $\mathbf{A}$ consists of all zeros.
For the HORIZONTAL-Coop scenario, each UE can connect to a certain number, say $\widetilde{N}_{\rm BS}$, of BSs or to a certain number, say $\widetilde{N}_{\rm AP}$, of APs. The association rule has to take into account the strength of the large-scale fading coefficients and the number of antennas at the APs and BSs. A suitable association rule can be obtained as follows:
@@
    then the $k$-th UE is to be associated to the APs with index $O^{\beta}_{k}(1),O^{\beta}_{k}(2),\ldots,O^{\beta}_{k}(\widetilde{N}_{\rm AP})$. Consequently, the coefficients $a_{k,O^{\beta}_{k}(1)},\ldots,a_{k,O^{\beta}_{k}(\widetilde{N}_{\rm AP})}$ should be set equal to 1, and all other entries in the matrix $\mathbf{A}$ and the entire matrix $\mathbf{B}$ should be set to zero.
@@
    If, on the contrary, (14) does not hold, then the $k$-th UE is to be associated to the BSs with index $O^{\rho}_{k}(1),O^{\rho}_{k}(2),\ldots,O^{\rho}_{k}(\widetilde{N}_{\rm BS})$. Consequently, the coefficients $b_{k,O^{\rho}_{k}(1)},\ldots,b_{k,O^{\rho}_{k}(\widetilde{N}_{\rm BS})}$ should be set equal to 1, and all other entries in the matrix $\mathbf{B}$ and the entire matrix $\mathbf{A}$ should be set to zero.
For the HET-noCoop scenario, where each UE can connect to either one BS or to one AP, the association rule UE-AP/BS is a special case of the one outlined for the above HORIZONTAL-Coop scenario, obtained by setting $\widetilde{N}_{\rm BS}=\widetilde{N}_{\rm AP}=1$.
Finally, in the FULL-Coop scenario, each UE is cooperatively served by $\widetilde{N}_{\rm AP}$ APs and by $\widetilde{N}_{\rm BS}$ BSs. Also for this scenario the large-scale fading coefficients are considered for selecting the serving APs and BSs. In particular, using again the sorting operators $O_{k}^{\beta}$ and $O_{k}^{\rho}$, we have that the association variables $a_{k,O^{\beta}_{k}(1)},\ldots,a_{k,O^{\beta}_{k}(\widetilde{N}_{\rm AP})}$ and $b_{k,O^{\rho}_{k}(1)},\ldots,b_{k,O^{\rho}_{k}(\widetilde{N}_{\rm BS})}$ are set equal to one, while the remaining entries of the matrices $\mathbf{A}$ and $\mathbf{B}$ are set to zero [^1] The cooperation between APs and/or BSs can be affected by timing synchronization errors whose consideration is beyond the scope of this paper. However, it is worth noting that, if every AP or BS is synchronized with its neighbors and the UE synchronizes with its Master AP or Master BS in the initial access, the model used in the paper can be assumed reasonable for analysis[9]. .
The above described association rules are, as already highlighted, simple and heuristic. Thus, alternative and more sophisticated rules can be conceived. In particular, the association could take into account not only the channel strength, as the described policies do, but also the interference level. Moreover, a load balancing constraint could also be introduced, in order to prevent the case in which a single AP or BS is connected to a too large number of devices. The final part of this paper will provide results along these lines.
@@ 4.2 Downlink beamforming
Once the UE-AP/BS association rule has been implemented, the set $\mathcal{K}_{m}$ ($\mathcal{J}_{\ell}$) can be defined for the generic $m$-th AP ($\ell$-th BS). In particular, $\mathcal{K}_{m}$ is the set of users served by the $m$th AP, that is, $\mathcal{K}_{m}=\{k\,:\,a_{k,m}=1\}$. Similarly, $\mathcal{J}_{\ell}$ is the set of users served by the $\ell$-th BS, that is, $\mathcal{J}_{\ell}=\{k\,:\,b_{k,\ell}=1\}$.
In this paper, the following three downlink beamformers will be considered.
@@
Partial zero-forcing (PZF): In this case, we assume that each AP performs a local PZF downlink precoding, i.e. it nulls its interference contribution towards some selected users. Consider the $m$-th AP and assume that $k\in\mathcal{K}_{m}$. Denote by $N_{\rm PZF}^{\rm AP}$ the number of users to be protected by the interference generated by each AP [^2] Recall that $N_{\rm PZF}^{\rm AP}$ must not exceed the number of antennas available at the AP. when transmitting to UE $k$. Denote by $j_{k,m}(1),\ldots,j_{k,m}(N_{\rm PZF}^{\rm AP})$ the index of the users to be protected [^3] A customary choice may be to select the users, other than the $k$-th, with the strongest large-scale fading coefficient towards the $m$-th AP.; the PZF precoder to be used at $m$-th AP to transmit to the $k$-th user is computed as follows:
@@
    Let $\widetilde{\mathbf{I}}_{k,m}=\mbox{orth}\left[\mathbf{I}_{k,m}\right]$ be a matrix that contains in its columns an orthonormal basis for the column span of $\mathbf{I}_{k,m}$.
@@
A similar procedure is used to determine the beamformer in $\ell$-th BS for transmission to UE $k$. Denote by $N_{\rm PZF}^{\rm BS}<N_{\rm BS}$ the number of users to be protected by the interference generated by BS $\ell$ when transmitting to UE $k$. Denote by $j_{k,\ell}(1),\ldots,j_{k,\ell}(N_{\rm PZF}^{\rm BS})$ the index of the users to be protected [^4] Again, these can be the users, other than the $k$-th, with the strongest large-scale fading coefficient towards the $\ell$-th BS.; the PZF precoder to be used at the $\ell$-th BS to transmit to $k$-th user thus requires the formation of the interference matrix $\mathbf{I}_{k,\ell}=[\widehat{\mathbf{h}}_{j_{k,\ell}(1),\ell},\ldots,\widehat{\mathbf{h}}_{j_{k,\ell}(N_{\rm PZF}^{\rm BS}),\ell}]$, containing the estimated channel of the users to be protected, the computation of an orthonormal basis $\widetilde{\mathbf{I}}_{k,\ell}=\mbox{orth}\left[\mathbf{I}_{k,\ell}\right]$, and of the beamformer: Define the PZF
@@
Minimum mean squares error (MMSE): The MMSE beamforming vector is obtained exploiting the duality principle in massive MIMO systems and minimizing the conditional mean-squares error (MSE) in the uplink decoding[34]. Assuming the LMMSE channel estimation discussed in Section 2.2.3, the MMSE beamforming vectors computed at the $m$-th AP and at the $\ell$-th BS for the $\mathcal{K}_{m}(i)$-th user and the $\mathcal{J}_{\ell}(i)$-th user are reported at the top of next page in Eqs. (16) and (17), respectively.
$\mathbf{w}_{\mathcal{K}_{m}(i),m}^{(a),{\rm MMSE}}=\frac{\left\{\displaystyle\sum_{j\in\mathcal{K}_{m}}\eta_{j}\left[\widehat{\mathbf{g}}_{j,m}\widehat{\mathbf{g}}_{j,m}^{H}+\mathbf{G}_{j,m}-\widehat{\mathbf{G}}_{j,m}\right]+\sigma^{2}_{z}\mathbf{I}_{N_{\rm AP}}\right\}^{-1}\widehat{\mathbf{g}}_{\mathcal{K}_{m}(i),m}}{\left\lVert\left\{\displaystyle\sum_{j\in\mathcal{K}_{m}}\eta_{j}\left[\widehat{\mathbf{g}}_{j,m}\widehat{\mathbf{g}}_{j,m}^{H}+\mathbf{G}_{j,m}-\widehat{\mathbf{G}}_{j,m}\right]+\sigma^{2}_{z}\mathbf{I}_{N_{\rm AP}}\right\}^{-1}\widehat{\mathbf{g}}_{\mathcal{K}_{m}(i),m}\right\rVert}\,,$ (16)
$\mathbf{w}_{\mathcal{J}_{\ell}(i),\ell}^{(b),{\rm MMSE}}=\frac{\left\{\displaystyle\sum_{j\in\mathcal{J}_{\ell}}\eta_{j}\left[\widehat{\mathbf{h}}_{j,\ell}\widehat{\mathbf{h}}_{j,\ell}^{H}+\mathbf{H}_{j,\ell}-\widehat{\mathbf{H}}_{j,\ell}\right]+\sigma^{2}_{z}\mathbf{I}_{N_{\rm BS}}\right\}^{-1}\widehat{\mathbf{h}}_{\mathcal{J}_{\ell}(i),\ell}}{\left\lVert\left\{\displaystyle\sum_{j\in\mathcal{J}_{\ell}}\eta_{j}\left[\widehat{\mathbf{h}}_{j,\ell}\widehat{\mathbf{h}}_{j,\ell}^{H}+\mathbf{H}_{j,\ell}-\widehat{\mathbf{H}}_{j,\ell}\right]+\sigma^{2}_{z}\mathbf{I}_{N_{\rm BS}}\right\}^{-1}\widehat{\mathbf{h}}_{\mathcal{J}_{\ell}(i),\ell}\right\rVert}\,,$ (17)